# AIDOVECL: AI-generated Dataset of Outpainted Vehicles for Eye-level Classification and Localization

### v1

@@ 3.2 Synthetic Data Generation
Seed images are randomly scaled and positioned on a $512\times 512512 × 512$-pixel blank canvas. The channels of these seed images are also permuted to diversify the colorization of vehicles. However, this does not alter black and white colors, which is partly advantageous as it maintains the black color of tires. Corresponding annotations are calculated based on the dimensions of the original bounding box (i.e., dividing the buffered seed dimensions by 1.15 to remove the 15% buffer), and the scale and position of the seed on the canvas. We also create a mask image to aid the outpainting modules. This mask maintains the buffer zone—unlike the annotation which accounts for the actual bounding box of the vehicle—and will be blurred on borders to ensure a smooth transition into the outpainted areas.
@@
For the outpainting process, we employ the stable diffusion inpainting model from Hugging Face (Paine, 2024a) which is based on LDM (Rombach et al., 2022). The model utilizes both positive and negative prompts. The positive prompt is constructed dynamically according to the vehicle class and follows the pattern: "A {location} during {time} with no vehicle." The {location} parameter can take values from {highway, road, street, downtown, plaza}, tailored to the vehicle type. For instance, TRUCK might use {highway, road, street}, while BUS is specifically linked to {street, downtown, plaza}. The {time} variable encompasses {spring, summer, fall, winter, a sunny day, a cloudy day, a rainy day, evening, sunset, sunrise}, providing a temporal context to complement the location. This structured prompt design ensures a contextually appropriate generation of the scene, enhancing the realism of the outpainted areas. To avoid introducing unwanted vehicles, which are not annotated in the previous step, the negative prompt explicitly excludes {traffic, train, car, truck, bus, van}. This selection relies on the fact that the masked area remains largely intact during outpainting, allowing us to maintain focus on our vehicle of the desired class. Additionally, we may add {billboard, text, advertisement} to the negative prompt to avoid generating images with billboards including brands which are on the vehicles and to minimize generating advertisement-like images of vehicles.
@@
We employ outpainting techniques to extend the boundaries of original vehicle images, creating visually plausible extended scenes. The quality of these outpainted images is rigorously evaluated using three no-reference image quality assessment scores: BRISQUE (Mittal et al., 2012), CLIP-IQA (Wang et al., 2023), and total variation (TV) loss (Rudin et al., 1992). BRISQUE is designed to assess natural scene statistics and quantify potential distortions, typically ranging from 0 to 100, with lower scores indicating better quality. Similarly, CLIP-IQA leverages the CLIP model’s ability to assess perceptual quality through deep learning, with scores between 0 and 1 where higher scores suggest superior perceptual quality. Furthermore, we calculate the TV loss for downscaled 32x32 images, a measure that promotes spatial smoothness by minimizing the intensity variation between adjacent pixels. This step is crucial for reducing the likelihood of selecting collage-like images. Utilizing low resolution for this calculation efficiently highlights global discontinuities often found in collage images. This makes it a practical approach for identifying abrupt changes that compromise the continuity of outpainted scenes. We set three criteria: a BRISQUE score of 15 or lower, a CLIP-IQA score of 0.9 or higher, and a maximum threshold of 15 for TV loss. Images that do not meet these thresholds are excluded from further consideration, ensuring that only high-quality, visually realistic extensions are pursued.
@@
To generate background images for our dataset, we utilize the text-to-image stable diffusion model from Hugging Face (Paine, 2024b) based on LDM (Rombach et al., 2022). This process involves selecting from predefined descriptions that vividly depict various urban scenes devoid of any vehicles, such as empty streets and parks at different times and under various atmospheric conditions. These textual prompts guide the generation of detailed and realistic background images while using negative prompts to ensure no vehicles appear. These background images are crucial for improving the performance of visual recognition algorithms by not only reducing false positives but also enhancing true negative rates. By providing clear and unobstructed scenes, the algorithms can more accurately learn what constitutes the absence of target objects, thereby improving overall detection reliability in complex urban environments.
@@ 3.3 Vehicle Classification and Localization
In this study, we utilize the YOLO (You Only Look Once) object detection framework (Redmon et al., 2016), specifically employing the YOLOv8 model (Ultralytics, 2024). We begin by loading a pretrained YOLOv8 model, which provides a robust starting point due to its weights being trained on a comprehensive dataset. The training process is conducted over 1000 epochs, using a batch size of 32, an initial learning rate of 0.01, and an image size of 512x512 pixels. An early stopping patience of 20 is set to prevent overfitting to the training data. For the augmented dataset, which is double the size, we double the batch size to 64 and initial learning rate to 0.02, maintaining the same number of model weight updates and ensuring effective learning progress. This approach leverages the efficiency and accuracy of pretrained models while adapting them to the unique characteristics of our dataset, assessing the object detection capabilities of AIDOVECL as a method to augment real vehicle images.
@@
In this work, we compare the performance of trained models using several key metrics: Precision, Recall, mAP50, mAP50-95, Fitness, and F1 Score. Precision measures the accuracy of positive predictions, indicating the proportion of correct positive identifications out of all positive predictions made. Recall evaluates the model’s ability to identify all relevant instances, representing the fraction of true positives detected among the actual positives. The mAP50 evaluates the model’s accuracy at a 50% Intersection over Union (IoU) threshold, assessing how well the predicted bounding boxes align with the ground truth. The mAP50-95 extends this evaluation across IoU thresholds from 50% to 95%, providing a robust measure of model performance under varying levels of detection stringency. Fitness is specifically defined in YOLOv8 and calculated as a weighted sum of metrics: 10% from mAP@0.5 and 90% from mAP@0.5:0.95, omitting weights for Precision and Recall by default. Finally, the F1 Score combines Precision and Recall into a single metric that quantifies the model’s overall accuracy and completeness in detection tasks.
@@ 3.4 Compute Resources
Our experiments were conducted on the Delta advanced computing and data resource at the National Center for Supercomputing Application (NCSA). The specific node utilized features 4 CPU cores, each allocated with 16 GB of memory, totaling 64 GB, and is equipped with an NVIDIA A100-SXM4-40GB GPU. The compute time for outpainting each $512\times 512512 × 512$-pixel image is approximately five seconds, and generating an image satisfying quality scores requires about 30 seconds on average.
@@ 4.1 Outpainted Vehicles
Samples of outpainted vehicles, as generated using the methodology described in the previous section, are illustrated in Figure 2. The diverse set of vehicle images results from varying the prompts for location and time of the scene, alongside the inherent randomness in generative inpainting models, and the seed’s color, scaling, and positioning on the canvas. Additionally, employing no-reference image quality assessments, such as TV loss, BRISQUE, and CLIP-IQA scores, ensures a degree of realism and natural appearance within the generated images. During attempts per every seed image, the method finds an outpainted image whose location and timing is more relevant to the type and lighting of the vehicle by meeting the quality criteria.
The scene integrity accross outpainted images underscores the capability of the inpainting model to generate relevant context around a vehicle, viewed at street level or generally from an eye-level perspective. Notably, the use of negative prompts to exclude unwanted vehicles has proven successful in almost all cases. This approach suggests that explicitly mentioning the masked object as a positive prompt is unnecessary, as the entire image generation is somehow conditioned upon it. Although it might seem counterintuitive, given that guidelines for inpainting models usually suggest including the masked object in the positive prompt, this approach helps us avoid background vehicles that are not self-annotated.

### v2

@@ 3.2 Synthetic Data Generation
Seed images are randomly scaled and positioned on a $512\times 512$-pixel blank canvas. The channels of these seed images are also permuted to diversify the colorization of vehicles. However, this does not alter black and white colors, which is partly advantageous as it maintains the black color of tires. Corresponding annotations are calculated based on the dimensions of the original bounding box (i.e., dividing the buffered seed dimensions by 1.15 to remove the 15% buffer), and the scale and position of the seed on the canvas. We also create a mask image to aid the outpainting modules. This mask maintains the buffer zone—unlike the annotation which accounts for the actual bounding box of the vehicle—and will be blurred on borders to ensure a smooth transition into the outpainted areas.
@@
For the outpainting process, we employ the stable diffusion inpainting model from Hugging Face (stable_diffusion_inpainting) which is based on LDM (rombach2022high). The model utilizes both positive and negative prompts. The positive prompt is constructed dynamically according to the vehicle class and follows the pattern: "A {location} during {time} with no vehicle." The {location} parameter can take values from { highway, road, street, downtown, driveway}, tailored to the vehicle type. The {time} variable encompasses { a sunny day, a cloudy day, evening, sunset, sunrise, dusk}, providing a temporal context to complement the location. This structured prompt design ensures a contextually appropriate generation of the scene, enhancing the realism of the outpainted areas. To avoid introducing unwanted vehicles, which are not annotated in the previous step, the negative prompt explicitly excludes {traffic, train, car, truck, bus, van}. This selection relies on the fact that the masked area remains largely intact during outpainting, allowing us to maintain focus on our vehicle of the desired class. Additionally, we may add {billboard, text} to the negative prompt to avoid generating images with billboards including brands which are on the vehicles and to minimize generating advertisement-like images of vehicles. Additional terms may be added to the negative prompt to suppress artifacts that compromise the realism and quality of the generated images.
@@
We employ outpainting techniques to extend the boundaries of original vehicle images, creating visually plausible extended scenes. The quality of these outpainted images is rigorously evaluated using three no-reference image quality assessment scores: BRISQUE (mittal2012no), CLIP-IQA (wang2023exploring), and QualiCLIP (agnolucci2024quality). BRISQUE is designed to assess natural scene statistics and quantify potential distortions, typically ranging from 0 to 100, with lower scores indicating better quality. Similarly, CLIP-IQA leverages the CLIP model’s ability to assess perceptual quality through deep learning, with scores between 0 and 1 where higher scores suggest superior perceptual quality. QualiCLIP is a self-supervised, opinion-unaware image quality assessment method that fine-tunes the CLIP model to align image representations with quality-related semantics through textual prompts, outputting a score between 0 and 1 where higher values indicate better perceptual quality. We retained images with a maximum BRISQUE score of 20, a minimum CLIP-IQA score of 0.5, and a minimum QualiCLIP score of 0.85, the latter serving as a high semantic quality bar to ensure that only images with strong alignment to high-quality visual semantics were included.
@@
To generate background images for our dataset, we utilize the text-to-image stable diffusion model from Hugging Face (stable_diffusion_v1_5) based on LDM (rombach2022high). This process involves selecting from predefined descriptions that vividly depict various urban scenes devoid of any vehicles, such as empty streets and parks at different times and under various atmospheric conditions. These textual prompts guide the generation of detailed and realistic background images while using negative prompts to ensure no vehicles appear. Background images are crucial for improving the performance of visual recognition algorithms by reducing false positives and enhancing the detection of true negatives. By providing clear and unobstructed scenes, the algorithms can more accurately learn what constitutes the absence of target objects, thereby improving overall detection reliability in complex urban environments.
Pipeline Parameters.
We encourage readers to consult the configuration file provided in the pipeline’s repository for the complete list of parameters used throughout the dataset generation process. These parameters—covering both seed image extraction, quality assessment, and outpainting—are primarily determined through iterative experimentation and empirical tuning, and are best understood in the context of the full pipeline setup.
@@ 3.3 Vehicle Classification and Localization
In this study, we utilize the YOLOv8 (You Only Look Once) object detection framework (redmon2016you), specifically employing the YOLOv8 model (ultralytics_docs) along with four other previously mentioned models (FCOS, RetinaNet, SSD, and FasterRCNN). We begin by loading pretrained models, which provide a robust starting point due to their weights being trained on a comprehensive dataset. The training process is conducted over 1000 epochs with an early stopping patience of 20 to prevent overfitting to the training data, using a batch size of 4. The learning rates were fixed at $2\times 10^{-3}$ for FCOS, $4\times 10^{-3}$ for RetinaNet, $1\times 10^{-3}$ for SSD, and $8\times 10^{-3}$ for Faster R-CNN, while YOLOv8 used Ultralytics’ default learning-rate configuration—including a starting rate of $0.01$, final fraction of $0.01$, warm-up over 3 epochs with momentum ramping from $0.8$ and bias learning rate of $0.1$, SGD momentum of $0.937$, weight decay of $0.0005$, and a linear learning rate decay schedule with cosine one-cycle scheduling disabled by default.
@@
In this work, we compare the performance of trained models using several key metrics: Precision, Recall, mAP50, mAP50-95, Fitness, and F1 Score. Precision measures the accuracy of positive predictions, indicating the proportion of correct positive identifications out of all positive predictions made. Recall evaluates the model’s ability to identify all relevant instances, representing the fraction of true positives detected among the actual positives. The mAP50 evaluates the model’s accuracy at a 50% IoU threshold, assessing how well the predicted bounding boxes align with the ground truth. The mAP50-95 extends this evaluation across IoU thresholds from 50% to 95%, providing a robust measure of model performance under varying levels of detection stringency. Fitness is specifically defined as a weighted sum of metrics: 10% from mAP@0.5 and 90% from mAP@0.5:0.95, omitting weights for Precision and Recall by default. Finally, the F1 Score combines Precision and Recall into a single metric that quantifies the model’s overall accuracy and completeness in detection tasks. In addition to overall metrics, we evaluate class-wise detection performance using confusion matrices and qualitative analysis of confused, false negative, and false positive instances.
@@ 3.4 Compute Resources
Our experiments were conducted on the Delta advanced computing and data resource at the National Center for Supercomputing Applications (NCSA). The specific node utilized features 4 CPU cores, each allocated with 16 GB of memory, totaling 64 GB, and is equipped with an NVIDIA A100-SXM4-40GB GPU. The compute time for outpainting each $512\times 512$-pixel image is approximately five seconds, and generating an image satisfying quality scores requires $30\sim 60$ seconds on average.
@@ 4 Results and Discussion
4.1 Comparing Detection Models for Seed Image Extraction
The vehicle seed images, designated for outpainting, are identified from a manually classified dataset of over 11,000 images using established object detection models. The selection of models for this task is determined by a consensus voting mechanism that considers pairs of models with an IoU exceeding 0.95, as previously described. Figure 2 displays the consensus level of each model, calculated as the percentage of affirmative votes received from other models, averaged across all vehicle images. For example, a consensus level of 0.5 indicates that the model agrees with two out of four possible models on average. As shown in the figure, FCOS receives the most approval, followed by RetinaNet, SSD, MaskRCNN, and FasterRCNN, reflecting their relative consensus rankings. We prioritize models for seed extraction in the specified order, using subsequent models as backups in the rare case that a model fails to detect an object.
Although this comparative study initially aimed to identify the most approved model within our pipeline, the resulting radar chart offers key insights for advancing future object detection models. For vehicle detection, the focus of this study, SSD, MaskRCNN, and FasterRCNN are less effective in bounding box extraction compared to FCOS and RetinaNet. Moreover, all models show a notably lower consensus level for trucks relative to other vehicle classes. This disparity is likely due to several reasons: frequent model failures in detecting trucks, possibly stemming from inadequate training set diversity. Additionally, the subjective nature of manual truck annotations in training data, where the trailer part might be inconsistently included, and the presence of advertisements on trucks, which complicates detection during inference, are significant challenges. To mitigate these issues, future detection models must be trained on more diverse datasets and ensure uniform and comprehensive annotation practices, particularly for complex vehicle categories like trucks.
Figure 2: Consensus level of every detection model is defined as the percentage of affirmative votes received from other models, averaged across all objects in that class.
@@ 4.2 Outpainted Vehicles
Samples of outpainted vehicles, as generated using the methodology described in the previous section, are illustrated in Figure 3. The diverse set of vehicle images results from varying the prompts for location and time of the scene, alongside the inherent randomness in generative inpainting models, and the seed’s scaling and positioning on the canvas. Additionally, employing no-reference image quality assessments, such as QualiCLIP, BRISQUE, and CLIP-IQA scores, ensures a degree of realism and natural appearance within the generated images. During attempts per every seed image, the method finds an outpainted image whose context is more relevant to the type and lighting of the vehicle by meeting the quality criteria.
The scene integrity across outpainted images underscores the capability of the inpainting model to generate relevant context around a vehicle, viewed at street level or generally from an eye-level perspective. Notably, the use of negative prompts to exclude unwanted vehicles has proven successful in almost all cases. This approach suggests that explicitly mentioning the masked object as a positive prompt is unnecessary, as the entire image generation is conditioned upon it. Although it might seem counterintuitive, this approach helps us avoid background vehicles that are not annotated automatically.
Figure 3: Outpainted images of various vehicle classes satisfying criteria for BRISQUE, CLIP-IQA, and QualiCLIP.